# Dynamic modes of active Potts models with factorizable numbers of states

### v1

@@ III.3 Four-State Potts Model
We first describe phase behavior of the four-state Potts model at thermal equilibrium. The four-state Potts model corresponds to the standard (two-color) Ashkin-Teller model, in which two Ising spins, $\sigma^{\mathrm{A}}=\pm 1$ and $\sigma^{\mathrm{B}}=\pm 1$, exist at each lattice site. The four types of pair states, i.e., $(\sigma^{\mathrm{A}},\sigma^{\mathrm{B}})=(1,1),(-1,1),(-1,-1)$, and $(1,-1)$, correspond to the four states $s=0$, $1$, $2$, and $3$ of the Potts model, respectively. The phase behavior at thermal equilibrium has been investigated in detail [21, 23, 24]. Under the conditions $J_{k,k}>J_{k,[k+1]}$ and $J_{k,k}>J_{k,[k+2]}$, three phases exist: a disordered phase with $\langle\sigma^{\mathrm{A}}\rangle=\langle\sigma^{\mathrm{B}}\rangle=\langle\sigma^{\mathrm{A}}\sigma^{\mathrm{B}}\rangle=0$, a Baxter (paramagnetic) ordered phase with $\langle\sigma^{\mathrm{A}}\rangle\neq 0$, $\langle\sigma^{\mathrm{B}}\rangle\neq 0$, and $\langle\sigma^{\mathrm{A}}\sigma^{\mathrm{B}}\rangle\neq 0$, and a partially ordered phase with $\langle\sigma^{\mathrm{A}}\rangle=\langle\sigma^{\mathrm{B}}\rangle=0$ and $\langle\sigma^{\mathrm{A}}\sigma^{\mathrm{B}}\rangle\neq 0$. In the partially ordered phase, two diagonal states ($s=0$ and $2$ or $s=1$ and $3$) are spatially mixed in the view of the Potts model, which corresponds to the M2 phase for the factorized number of states ($4=2\times 2$) in the present study.
Figure 18: Four-state Potts model at $L=128$. (a–d) Snapshots. The light-yellow, light-green, medium-blue, and red sites (light to dark in grayscale) represent $s=0$, $1$, $2$, and $3$, respectively. (a) Nonspiral waves (W4) at $h=1.1$ and $J_{k,[k+2]}=0$. (b) Spiral waves (W4) at $h=1$ and $J_{k,[k+2]}=-2$. (c) Homogeneous cycling of two diagonal states (HC2) at $h=1$ and $J_{k,[k+2]}=1$. (d) Mixing of two diagonal states (M2) at $h=1$ and $J_{k,[k+2]}=1.5$. (e) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, and green down-pointing triangles represent HC4, HC2, M2, W4, and WI, respectively.
@@
Here, under active flipping of $h>0$, the M2 phase persists when $J_{k,[k+2]}\gtrsim 1$ (see Fig. 18(d) and (e)). For intermediate values of $J_{k,[k+2]}$, homogeneous cycling (switching) of two diagonal phases (HC2) emerges (see Fig. 18(c) and (e)). In the $k$-state dominant phase, the nuclei of $[k+1]$ state are swallowed up by the $[k+2]$ domains. In the M2 mode, no switching between even- and odd-numbered states occur in the considered simulation periods. These behaviors are similar to those of the M3 and HC3 modes of the six-state models. The transition between the HC2 and M2 modes is continuous, and the transition point is detected based on $U_{1}$ (see Fig. 19(b)). Since $R_{1}=0.57\pm 0.01$ at the transition point, we consider the system is taken as being in the M2 mode when $R_{1}<0.57$.
@@ III.4 Eight-State Potts Model
Figure 20: Eight-state Potts model at $J_{k,[k+3]}=J_{k,[k+4]}=J_{k,[k+2]}$ and $L=128$. (a) Snapshot of waves of eight states (W8) at $h=1.3$ and $J_{k,[k+2]}=0$. (b) Snapshot of waves of four states (W4) at $h=1$ and $J_{k,[k+2]}=0.6$. The white, yellow, green, light-blue, dark-blue, magenta, red, and black sites (light to dark in grayscale) represent $s=0$, $1$, $2$, $3$, $4$, $5$, $6$, and $7$, respectively. (c) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, and green down-pointing triangles represent HC8, W4, M8, W8, and WI, respectively.
Figure 21: Dependence on $J_{k,[k+2]}$ in the eight-state Potts model at $J_{k,[k+3]}=J_{k,[k+4]}=J_{k,[k+2]}$ and $h=1.5$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for $k$$[k+1]$, $k$$[k+2]$, $k$$[k+3]$, and $k$$[k+4]$ pairs. (c) Mean cluster size $n_{\mathrm{c1}}$. The solid lines represent the data at $L=128$. In (b) and (c), the dashed lines with crosses represent the data at $L=64$.
@@ IV Relation with Other Models
where $J,K>0$ and $\epsilon_{\alpha\beta}$ is the Levi-Civita symbol. The energy of the first term decreases when the spins of the same type in the neighboring sites are in paramagnetic order. However, the second term gives a nonreciprocal intra-site interaction, such that the spin A is preferentially antimagnetic to the spin B at the same site ($\sigma_{i}^{A}\sigma_{i}^{B}<0$), whereas the spin B is preferentially paramagnetic ($\sigma_{i}^{A}\sigma_{i}^{B}>0$). Thus, the following cyclic change is induced: $(\sigma^{\mathrm{A}},\sigma^{\mathrm{B}})=(1,1)\to(-1,1)\to(-1,-1)\to(1,-1)\to(1,1)$. When the Avni model is mapped to the present active Potts model, $h=2K$, $J_{k,k}=2J$, $J_{k,[k+1]}=0$, and $J_{k,[k+2]}=-2J$ are obtained. The contact energy is expressed as $J_{k,[k+j]}=2J\cos(\pi j/2)$, such that these interactions correspond to those of the four-state vector Potts model (also called clock model) [16]. In the phase diagram of Fig. 18(e), the data obtained at $J_{k,[k+2]}=-2$ correspond to those of the Avni model at $J=1$. At low and high $h$, the HC4 and W4 modes occur, respectively, and the two modes temporally coexist around the transition point in the long-term dynamics.
Lattice Lotka–Volterra models, based on predator–prey interactions, have been widely used to study spatiotemporal patterns [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Cyclic changes in species (states) are considered similar to those of the present models. The Lotka–Volterra models of four or more species can exhibit short-term dynamics similar to those given by the present model, such as the long-lived spatial coexistence of two phases and domains of mixed species [41, 46, 47, 48, 49, 50, 51, 52, 53, 54]. However, the long-term dynamics completely differ. In the long-term limit, the Lotka–Volterra models exhibit only a uniform phase occupied by a single species through an absorbing transition. Extinct species never reappear, because species multiply through self-reproduction. Technically, this is proceeded by pair updates, such as $(s_{i},s_{j})=(0,1)\to(1,1)$, in which the $s_{i}$ species becomes the $s_{j}$ species, indicating predation and self-reproduction. These updates involve no backward processes; therefore, a detailed balance cannot be obtained. We compared the single update with this pair update using a three-state Potts model and demonstrated that the pair update generates fewer spirals at high flipping energy, but all sites are occupied by a single state in the long-term limit under all conditions (including $h=J_{ss^{\prime}}=0$[34]. Therefore, an update scheme including backward processes is crucial to obtain long-term spatiotemporal dynamics.
@@ V Summary
We studied the nonequilibrium dynamics of the $q$-state active Potts models with $q=4$, $5$, $6$, and $8$ under cyclically symmetric conditions in the long-term limit. Three mode types and their combinations were identified: mixed phases, wave propagation modes, and cycling of homogeneous phases. In addition to these modes comprising $q$ states, modes comprising a factorized number of states emerge for $q=4$, $6$, and $8$: a mixed phase and homogeneous switching of two diagonal states for $q=4$, a wave mode of four states for $q=8$, and several modes for $q=6$. In the six-state model, two types of spiral waves with factorized symmetry were found: spiral waves of three states (three odd numbers, $s=1,3$, and $5$, or three even numbers, $s=0,2$, and $4$), with the other three states present as small clusters at the domain boundaries; and spiral waves of three types of mixed domains, in which two diagonal states ($s=k$ and $[k+3]$) are mixed within a domain. In the former and latter spirals, one flip cycle in the spiral domains requires one cycle and half cycle of states, respectively. Moreover, homogeneous cycling modes of the three odd- or even-numbered states and three mixed phases were found. Therefore, the factorization of symmetry can generate rich dynamic modes under nonequilibrium conditions, including the combinations of static (mixing) and dynamic (wave) modes.
@@
We calculated the scaling exponents of the second-order transitions between the W3 and M6 modes and between the HC3 and M3 modes. The exponents of the W3–M6 transition are modified from the equilibrium values, whereas no change is detected for the HC3–M3 transition. The transitions to wave modes may generally exhibit greater changes in the exponents than those between static modes. In contrast, first-order transitions have not been found yet at $N\to\infty$ in the present models. Temporal coexistences of the HC and wave modes occur, but the transition points shift to $h=0$ with increasing system size. Since the forward and backward transitions typically take different pathways, they can exhibit different $N$ dependencies. Clarifying the condition for first-order transitions at $N\to\infty$ is an interesting problem for further studies.
@@
In this study, we have kept the cyclic symmetry and fixed the temperature and contact energies between the same states and between neighboring states. In our previous studies, we reported non-cyclic homogeneous phases and the coexistence of two diagonal phases under asymmetric conditions (for three- and four-state Potts models) [35, 36]. For $q\geq 5$, the imposition of asymmetric conditions will likely modify the dynamic modes in a similar manner; however, symmetry factorization may additionally induce different types of dynamics. Even restricting to the symmetric conditions, we have not explored the entire parameter space. Under equilibrium, several antimagnetic phases have been reported for the Ashkin-Teller models [21, 22, 23, 24]. Thus, antimagnetic and other mode types likely exist in the remaining parameter space in active Potts models. Further, the modes and transitions can be changed in three (or higher)-dimensional space. Overall, the active Potts models provide a simple but rich system for exploring various types of nonequilibrium dynamics under thermal fluctuations.

### v2

@@ III.3 Four-State Potts Model
We first describe phase behavior of the four-state Potts model at thermal equilibrium. The four-state Potts model corresponds to the standard (two-color) Ashkin-Teller model, in which two Ising spins, $\sigma^{\mathrm{A}}=\pm 1$ and $\sigma^{\mathrm{B}}=\pm 1$, exist at each lattice site. The four types of pair states, i.e., $(\sigma^{\mathrm{A}},\sigma^{\mathrm{B}})=(1,1),(-1,1),(-1,-1)$, and (1,1)(1,-1), correspond to the four states $s=0$, 11, 22, and $3$ of the Potts model, respectively. The phase behavior at thermal equilibrium has been investigated in detail [21, 23, 24]. Under the conditions $J_{k,k}>J_{k,[k+1]}$ and $J_{k,k}>J_{k,[k+2]}$, three phases exist: a disordered (paramagnetic) phase with $\langle\sigma^{\mathrm{A}}\rangle=\langle\sigma^{\mathrm{B}}\rangle=\langle\sigma^{\mathrm{A}}\sigma^{\mathrm{B}}\rangle=0$, a Baxter (ferromagnetic) ordered phase with $\langle\sigma^{\mathrm{A}}\rangle\neq 0$, $\langle\sigma^{\mathrm{B}}\rangle\neq 0$, and $\langle\sigma^{\mathrm{A}}\sigma^{\mathrm{B}}\rangle\neq 0$, and a partially ordered phase with $\langle\sigma^{\mathrm{A}}\rangle=\langle\sigma^{\mathrm{B}}\rangle=0$ and $\langle\sigma^{\mathrm{A}}\sigma^{\mathrm{B}}\rangle\neq 0$. In the partially ordered phase, two diagonal states ($s=0$ and $2$ or $s=1$ and $3$) are spatially mixed in the view of the Potts model, which corresponds to the M2 phase for the factorized number of states ($4=2\times 2$) in the present study.
Figure 18: Four-state Potts model at $L=128$. (a–d) Snapshots. The light yellow, green, blue, and red sites (light to dark in grayscale) represent $s=0$, 11, 22, and $3$, respectively. (a) Nonspiral waves (W4) at $h=1.1$ and $J_{k,[k+2]}=0$. (b) Spiral waves (W4) at $h=1$ and $J_{k,[k+2]}=-2$. (c) Homogeneous cycling of two diagonal states (HC2) at $h=1$ and $J_{k,[k+2]}=1$. (d) Mixing of two diagonal states (M2) at $h=1$ and $J_{k,[k+2]}=1.5$. (e) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, and green down-pointing triangles represent HC4, HC2, M2, W4, and WI, respectively.
@@
Here, under active flipping of $h>0$, the M2 phase persists when $J_{k,[k+2]}\gtrsim 1$ (see Fig. 18(d) and (e)). For intermediate values of $J_{k,[k+2]}$, homogeneous cycling (switching) of two diagonal phases (HC2) emerges (see Fig. 18(c) and (e)). In the $k$-state dominant phase, the nuclei of $[k+1]$ state are swallowed up by the $[k+2]$ domains. In the M2 mode, no switching between even- and odd-numbered states occur in the considered simulation periods. These behaviors are similar to those of the M3 and HC3 modes of the six-state models. The transition between the HC2 and M2 modes is continuous, and the transition point is detected based on $U_{1}$ (see Fig. 19(b)). Since $R_{1}=0.57\pm 0.01$ at the transition point for $h=1$, we consider the system is taken as being in the M2 mode when $R_{1}<0.57$.
@@
Figure 20: Eight-state Potts model at $J_{k,[k+3]}=J_{k,[k+4]}=J_{k,[k+2]}$ and $L=128$. (a) Snapshot of waves of eight states (W8) at $h=1.3$ and $J_{k,[k+2]}=0$. (b) Snapshot of waves of four states (W4) at $h=1$ and $J_{k,[k+2]}=0.6$. The white, yellow, light green, cyan, blue, magenta, red, and black sites (light to dark in grayscale) represent $s=0$, 11, 22, 33, 44, 55, 66, and $7$, respectively. (c) Dynamic phase diagram. The red circles, gray crosses, magenta diamonds, blue up-pointing triangles, and green down-pointing triangles represent HC8, W4, M8, W8, and WI, respectively.
@@ III.4 Eight-State Potts Model
In general, the factorized symmetry can emerge when the domain contact of factorized number of states is more stable than the other contacts. For $q=2n$, the contact of even or odd number of states ($s=2n^{\prime}$ or $2n^{\prime}+1$) can be stabilized by increasing $J_{k,[k+2]}$. Similarly, $s=3n^{\prime}$, $3n^{\prime}+1$, or $3n^{\prime}+2$ by increasing $J_{k,[k+3]}$ for $q=3n$. Therefore, the waves of factorized number of states are likely formed also for $q>8$.
Figure 21: Dependence on $J_{k,[k+2]}$ in the eight-state Potts model at $J_{k,[k+3]}=J_{k,[k+4]}=J_{k,[k+2]}$ and $h=1.5$. (a) Probabilities $p_{\mathrm{phase}}$ of one-phase and multi-phase coexistence states. The bidirectional arrows at the top represent the ranges of dynamic modes. (b) Contact probabilities $p_{\mathrm{contact}}$ for kk[k+1][k+1], kk[k+2][k+2], kk[k+3][k+3], and kk[k+4][k+4] pairs. (c) Mean cluster size $n_{\mathrm{c1}}$. The solid lines represent the data at $L=128$. In (b) and (c), the dashed lines with crosses represent the data at $L=64$.
@@ IV Relation with Other Models
where $J,K>0$ and $\epsilon_{\alpha\beta}$ is the Levi-Civita symbol. The energy of the first term decreases when the spins of the same type in the neighboring sites are in ferromagnetic order. However, the second term gives a nonreciprocal intra-site interaction, such that spin A is preferentially antiferromagnetic to spin B at the same site ($\sigma_{i}^{A}\sigma_{i}^{B}<0$), whereas spin B is preferentially ferromagnetic ($\sigma_{i}^{A}\sigma_{i}^{B}>0$). Thus, the following cyclic change is induced: $(\sigma^{\mathrm{A}},\sigma^{\mathrm{B}})=(1,1)\to(-1,1)\to(-1,-1)\to(1,-1)\to(1,1)$. When the Avni model is mapped to the present active Potts model, $h=2K$, $J_{k,k}=2J$, $J_{k,[k+1]}=0$, and $J_{k,[k+2]}=-2J$ are obtained. The contact energy is expressed as $J_{k,[k+j]}=2J\cos(\pi j/2)$, such that these interactions correspond to those of the four-state vector Potts model (also called the clock model) [16]. In the phase diagram of Fig. 18(e), the data obtained at $J_{k,[k+2]}=-2$ correspond to those of the Avni model at $J=1$. At low and high $h$, the HC4 and W4 modes occur, respectively, and the two modes temporally coexist around the transition point in the long-term dynamics.
Lattice Lotka–Volterra models, based on predator–prey interactions, have been widely used to study spatiotemporal patterns [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Cyclic changes in species (states) are considered similar to those of the present models. The Lotka–Volterra models of four or more species can exhibit short-term dynamics similar to those given by the present model, such as the long-lived spatial coexistence of two phases and domains of mixed species [41, 46, 47, 48, 49, 50, 51, 52, 53, 54]. However, the long-term dynamics completely differ. In the long-term limit, the Lotka–Volterra models exhibit only a uniform phase occupied by a single species through an absorbing transition. Extinct species never reappear, because species multiply through self-reproduction. Technically, this is proceeded by pair updates, such as $(s_{i},s_{j})=(0,1)\to(1,1)$, in which the $s_{i}$ species becomes the $s_{j}$ species, indicating predation and self-reproduction. These updates involve no backward processes; therefore, a detailed balance cannot be obtained. We compared the single update with this pair update using a three-state Potts model and demonstrated that the pair update generates fewer spirals at high flipping energy, but all sites are occupied by a single state in the long-term limit under all conditions (including $h=J_{ss^{\prime}}=0$[34]. In the case that flips between neighboring states are only allowed for $q\geq 4$, a pattern becomes frozen when the contact of neighboring states disappears. However, the absorbing transition to a uniform state occurs when flips between non-neighboring states with $h_{kk^{\prime}}=0$ are added ($k^{\prime}=[k+2]or[k+3]$ for $q=6$). Therefore, an update scheme including backward processes is crucial to obtain long-term spatiotemporal dynamics.
@@ V Summary
We studied the nonequilibrium dynamics of the $q$-state active Potts models with $q=4$, 55, 66, and $8$ under cyclically symmetric conditions in the long-term limit. Three mode types and their combinations were identified: mixed phases, wave propagation modes, and cycling of homogeneous phases. In addition to these modes comprising $q$ states, modes comprising a factorized number of states emerge for $q=4$, 66, and $8$: a mixed phase and homogeneous switching of two diagonal states for $q=4$, a wave mode of four states for $q=8$, and several modes for $q=6$. In the six-state model, two types of spiral waves with factorized symmetry were found: spiral waves of three states (three odd numbers, $s=1,3$, and $5$, or three even numbers, $s=0,2$, and $4$), with the other three states present as small clusters at the domain boundaries; and spiral waves of three types of mixed domains, in which two diagonal states ($s=k$ and $[k+3]$) are mixed within a domain. In the former and latter spirals, one flip cycle in the spiral domains requires one cycle and half cycle of states, respectively. Moreover, homogeneous cycling modes of the three odd- or even-numbered states and three mixed phases were found. Therefore, the factorization of symmetry can generate rich dynamic modes under nonequilibrium conditions, including the combinations of static (mixing) and dynamic (wave) modes.
@@
We calculated the scaling exponents of the second-order transitions between the W3 and M6 modes and between the HC3 and M3 modes. The exponents of the W3–M6 transition are modified from the equilibrium values, whereas no change is detected for the HC3–M3 transition. The transitions to wave modes may generally exhibit greater changes in the exponents than those between static modes. In contrast, first-order transitions have not been reported yet at $N\to\infty$ in the present models. Temporal coexistences of the HC and wave modes occur, but the transition points shift to $h=0$ with increasing system size. Since the forward and backward transitions typically take different pathways, they can exhibit different $N$ dependencies. However, first-order transitions can be obtained in asymmetric conditions [65].
@@
In this study, we have kept the cyclic symmetry and fixed the temperature and contact energies between the same states and between neighboring states. In our previous studies, we reported non-cyclic homogeneous phases and the coexistence of two diagonal phases under asymmetric conditions (for three- and four-state Potts models) [35, 36]. For $q\geq 5$, the imposition of asymmetric conditions will likely modify the dynamic modes in a similar manner; however, symmetry factorization may additionally induce different types of dynamics. Even restricting to the symmetric conditions, we have not explored the entire parameter space. Under equilibrium, several antiferromagnetic phases have been reported for the Ashkin-Teller models [21, 22, 23, 24]. Thus, antiferromagnetic and other mode types likely exist in the remaining parameter space in active Potts models. Further, the modes and transitions can be changed in three (or higher)-dimensional space. Overall, the active Potts models provide a simple but rich system for exploring various types of nonequilibrium dynamics under thermal fluctuations.